a spark plug in an aoutomobile engine consists of two metal conductors that are separated by a distance of 0.5mm. When an electric spark jumps between them, the magnitude of the electric field is 4.0x107V/m. What is the magnitude of the potential difference V between the conductors?
1
Expert's answer
2021-04-29T10:44:22-0400

Separation between plates, "d=0.5\space mm=0.5\times10^{-3}\space m"

Electric field, "E=4\times10^7\space V/m"

Electric field between two conducting plates, "E=\dfrac{V}{d}"

"V=E.d=(4\times10^7)\times(0.5\times10^{-3})=2\times10^4\space V"